OpenStudy (anonymous):

1. alicia and jamal completed the following " given : 3x+7=x-5 prove: x=-6 who completed the proof correctly?

OpenStudy (texaschic101):

I can't see how each one completed it...but I would do it this way.. 3x + 7 = x - 5 --- subtract x from each side 3x - x + 7 = -5 -- now subtract 7 from both sides 3x - x = -5 - 7 -- now simplify 2x = - 12 -- now divide both sides by 2 x = -6

OpenStudy (texaschic101):

but there are other ways

OpenStudy (anonymous):

this is how 3x+7=x-5 : given 2x+7=-5: subtraction property of equality 2x=-12: subtraction property of quality x=-6: division property of equality for alicia

OpenStudy (anonymous):

and jamals 3x+7=x-5 : given 7=-2x-5 : subtraction property of equality 12=-2x: addition property of equality -6=x : division property of equality

OpenStudy (anonymous):

a) alicia only b) jamal only c) both alicia and jamal d) neither alicia nor jamal

OpenStudy (texaschic101):

oh....they both look correct to me....glad you put the answer choices...I was thinking there could only be one of them right..lol....they are both correct

OpenStudy (anonymous):

thanks so much

OpenStudy (texaschic101):

sure thing :)